# Landau and leading singularities in arbitrary space-time dimensions

### v2

@@ Acknowledgments
We are indebted to Johannes Henn for discussions and feedback on the manuscript, and Janusz Gluza and Einan Gardi for comments on the manuscript. We also wish to thank Jungwon Lim, Sebastian Mizera, and Andrew McLeod for insightful discussions. We are thankful to the anonymous Referee for carefully reading our manuscript and for providing useful comments that led to an improved version of the results summarised in this communication. This research received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 725110), Novel structures in scattering amplitudes. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix B Supplementary results
$\frac{1}{D_{1}\ldots D_{m}D_{m+1}}$ $=\frac{1}{D_{m+1}}\sum_{i=1}^{m}\frac{1}{D_{i}}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}D_{ji}}$ (158)
@@
$\sum_{i=1}^{m+1}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m+1}D_{ji}}=$ (167)
$=\sum_{i=1}^{m}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m+1}D_{ji}}+\sum_{i=1}^{m}\frac{1}{D_{i,m+1}}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}\left(D_{j,m+1}-D_{i,m+1}\right)}$ (172)
@@ Appendix C Additional examples of one-loop Landau singularities
The results reported in Eqs. (IV.1) and (IV.1) were obtained through the use of the following identities,
@@
$\det(A)=\begin{vmatrix}2k_{1}&\ldots&-2k_{n}\\ 2p_{11}&\ldots&-2p_{1n}\\ \vdots&\ddots&\vdots\\ 2p_{(n-1)1}&\ldots&-2p_{(n-1)n}\end{vmatrix}=$ (207)

### v3

@@ Acknowledgments
We are indebted to Johannes Henn for discussions and feedback on the manuscript, and Janusz Gluza and Einan Gardi for comments on the manuscript. We also wish to thank Jungwon Lim, Sebastian Mizera, and Andrew McLeod for insightful discussions. We are thankful to the anonymous Referee for carefully reading our manuscript and for providing useful comments that led to an improved version of the results summarised in this communication. This research received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 725110), Novel structures in scattering amplitudes, the Polish National Science Center (NCN) under grant 2023/50/A/ST2/00224, and by the Leverhulme Trust, LIP-2021-01. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix B Supplementary results
$\frac{1}{D_{1}\ldots D_{m}D_{m+1}}$ $=\frac{1}{D_{m+1}}\sum_{i=1}^{m}\frac{1}{D_{i}}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}D_{ji}}$
@@
$\sum_{i=1}^{m+1}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m+1}D_{ji}}=$
$=\sum_{i=1}^{m}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m+1}D_{ji}}+\sum_{i=1}^{m}\frac{1}{D_{i,m+1}}\frac{1}{\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{m}\left(D_{j,m+1}-D_{i,m+1}\right)}$
@@ Appendix C Additional examples of one-loop Landau singularities
The results reported in Eqs. (54) and (58) were obtained through the use of the following identities,
@@
$\det(A)=\begin{vmatrix}2k_{1}&\ldots&-2k_{n}\\ 2p_{11}&\ldots&-2p_{1n}\\ \vdots&\ddots&\vdots\\ 2p_{(n-1)1}&\ldots&-2p_{(n-1)n}\end{vmatrix}=$
@@
Data Availability Statement:
No Data associated in the manuscript